3. The 100g block, which moves at a speed of 4m / s, starts to brake until it stops. The work done in this case is:

 A) -800 J

 B) 8 J

 C) 0.8J

 D) -0.8J


1
Expert's answer
2021-03-29T08:56:51-0400

"W=\Delta KE=mv^2/2-mv_0^2/2=0-0.1\cdot4^2/2=-0.8\ (J)" . Answer